OpenStudy (anonymous):

How many orbitals are in the 3s subshell?

OpenStudy (rogue):

How many orbital(s) does any s-subshell hold?

OpenStudy (anonymous):

0

OpenStudy (rogue):

:O no! The s-subshell has 1 orbital, so it can hold 2 electrons. p-subshell has 3 orbitals, and so it can hold 6 electrons. d-subshell has 5 orbitals and it can hold 10 e-. And it goes on and on...

OpenStudy (rogue):

That answers your next question =P

OpenStudy (anonymous):

I'm sorry, I was thinking about quantum number

OpenStudy (rogue):

s has an azimuthal quantum # of 0. The principle quantum number can be anything > 0 though.

OpenStudy (anonymous):

in that case, how many orbitals are there in the third shell, n=3

OpenStudy (rogue):

# of orbitals in a shell = n^2 # of electrons in a shell = 2 * n^2

OpenStudy (rogue):

Seems like everyone is starting quantum mechanics ;D

OpenStudy (anonymous):

so it would be 18

OpenStudy (anonymous):

lol i know right, it's killing me but I'm understanding it better now

OpenStudy (rogue):

9 orbitals, 18 e-

OpenStudy (anonymous):

Thank you so much

OpenStudy (rogue):

yw :)